OpenStudy (anonymous):

x^2 -6x-16=0 ( I need the steps for my answer this what I have for answer -2,8 but I dont know the steps

OpenStudy (anonymous):

(x+2)(x-8)=0 x+2=0 => x= -2 x-8=0 => x= 8

OpenStudy (anonymous):

x^2-6x-16=0 x^2+2x-8x-16=0 x(x+2)-8(x+2)=0 (x-8)(x+2)=0 then audia4 has displayed